OpenStudy (anonymous):

use multiplying by 1 to fine an exression equivalent to (9) (4) with a denominator of 4c

OpenStudy (anonymous):

(4c)/(4c) = 1 if c is non-zero. (9)*(4)*1 = (9)*(4)*(4*c)/(4*c) = 144c/(4c)